Zeronknight (zeronknight):

What is the opposite of this inequality? -x\leq -8 A. x\leq 8 B. x\geq -8 C. x\leq -8 D. x\geq 8

Zeronknight (zeronknight):

\[-x\leq -8\]

Zeronknight (zeronknight):

@mathmale @pooja195

OpenStudy (mathmale):

Question is hard to decipher. What do you mean by "opposite" in this context?

OpenStudy (mathmale):

-x\leq -8 is a bit hard to read. \[-x\leq -8\]

OpenStudy (mathmale):

is the result I get when I copy and paste -x\leq -8 into the Equation Editor screen.

OpenStudy (mathmale):

You might want to solve\[-x\leq -8\]

Zeronknight (zeronknight):

Hmmm it asks the opposite of that. I could do the answers as well fi you want XD

OpenStudy (mathmale):

... for x.

Zeronknight (zeronknight):

I don't understand how do I do the opposite? Do I add -8 +8?

OpenStudy (mathmale):

I don't understand what you mean by "opposite" in this context. I'd suggest you solve \[-x\leq -8\]

OpenStudy (mathmale):

... for x and then write a NEW SET for the part of the number line not described by \[-x\leq -8\]

Zeronknight (zeronknight):

ohhh

OpenStudy (mathmale):

To do this: 1) Mult. both sides of the given inequality by (-1) 2) Reverse the direction of the inequaltiy sign

Zeronknight (zeronknight):

how would I multiply both sides? And is negative times negative, positive?

OpenStudy (mathmale):

\[(-1)[-x\leq -8]\]

OpenStudy (mathmale):

Hope the above answers your question. And yes, neg times neg is pos.

OpenStudy (mathmale):

(-1)(-x)=?

OpenStudy (mathmale):

(-1)(-8)=?

Zeronknight (zeronknight):

-1 x -8 = 8. -1 x -x = x?

OpenStudy (mathmale):

Very important that you include parentheses here for clarity. See what I did, above. (-1)(-x) = x. that's it.

OpenStudy (mathmale):

After y ou've done this, change the direction of the inequality sign.

OpenStudy (mathmale):

I get \[x \ge 8.\]

Zeronknight (zeronknight):

ohhh

OpenStudy (mathmale):

"x is equal to or greater than 8." I'd suppose that the "opposite" to which you were referring would be the part of the number line NOT covered by "x is equal to or greater than 8". that would be "x is less than 8." I need to excuse myself right now, but will be back on OpenStudy in an hour or so.

OpenStudy (yumyum247):

D